Solve 3n2-6n+4=0 using completing the square method.

Answer by richwmiller(17219)

3n^2-6n+4=0
n^2-2n+4/3=0
n^2-2n=-4/3
(-2/2)^2=1
n^2-2n+1=-1/3
(n-1)^2=-1/3